suppose that the demand function or schedule is given by the linear equation p = 300 - qd and the supply equation by p =60 + 2 qs (a) find the equilibrium price and quantity (b) suppose the supply curve slop prices from 2 qs to 3 qs (C) find the new equilibrium price and quantity (d) find the equilibrium position graphically?




1
Expert's answer
2021-07-30T03:02:01-0400

"(a)\\
Qd=300-p\\
Qs=p-60\div2\\
Qd=Qs\\
300-p=p-60\div2\\
600-2p=p-60\\
660=3p\\
p=110\\
Qd=Qs=300-p\\
=300-110\\=190"


(b)

"p=60+3Qs\\
Qs=p-60\div3\\
Qd=300-p\\
Qs=Qd\\
p-60\div3=300-p\\
p-60=900-3p\\
4p=900+60\\
p=960\div4\\
p=240"